# Capturing Stable HDR Videos Using a Dual-Camera System

### v1

@@ IV-E Loss Function
$\mathcal{A}\mathcal{S}\mathcal{L}(\widehat{Z},Z)=\frac{1}{N}\sum\mid\operatorname{Sobel}^{\star}(\widehat{Z})-\text{ Sobel }^{\star}(Z)\mid\$ (13)
@@
$\operatorname{Loss}(\widehat{Z},Z)=\mathcal{L}1(\widehat{Z},Z)+\lambda\cdot\mathcal{D}-\mathcal{AS}\mathcal{L}^{D}(\widehat{Z},Z)$ (15)
@@ V EXPERIMENTS
V-1 Settings
@@
The proposed method is implemented in PyTorch on an NVIDIA RTX4090 GPU. For the convolutional layers, we employ 64 kernels of size 3$\times$3 with a stride of one, applying zero padding to preserve the dimensions of the resulting feature maps. The optimization is performed using the Adam optimizer [57] with an initial learning rate of $10^{-4}$, and the training concludes when the learning rate reaches $10^{-6}$. To train the model, both the LDR and corresponding images are cropped into 256 $\times$ 256 patches. A batch size of 16 is utilized during training. Patches are randomly rotated for data augmentation to avoid overfitting during the training stage.
V-2 Datasets
We employ four benchmark datasets for evaluation: Prabhakar’s dataset [30], Kalantari’s dataset [28], Tursun’s dataset [58] and Sen’s dataset [5]. Details of these datasets are described as follows:
@@ V-2 Datasets
1). Prabhakar’s Dataset. Including 582 varying exposure images with corresponding deghosted HDR images, with exposure values between -3EV to +3EV.
@@
2). Kalantari’s dataset. Including 74 training samples and 15 testing samples. Each consists of three LDR images, and the exposure values belong to {-2, 0, +2} or {-3, 0, +3}.
@@
3). Tursun’s dataset and Sen’s dataset. Lacking ground truth, is used to assess the generalization ability of the models.
V-3 Metrics
@@
We quantitatively evaluate the predicted HDR images by measuring PSNR and SSIM [59] in both the linear domain (-L) and the tone-mapped domain (-$\mu$) [28]. The tone mapping operation is described in Sec. IV-E. Additionally, we compare the results using HDR-VDP-2 [60] for conforming the human visual perception. We also quantitatively evaluated the luminance consistency of the video by testing the Luminance Standard Deviation (LSD), temporal SSIM (t-SSIM) and average luminance ($L_{avg}$):
@@
V-A Comparison with state-of-the-art methods
We quantitatively and qualitatively compare our EAFNet to several state-of-the-art MEF methods including Kalantari [49], AHDR [36], Prabhakar [61], HDR-Transformer [40], SCTNet [43], DomainPlus [39], SAFNet [31], and several state-of-the-art AE methods including HDRFlow [51], LAN-HDR [50], and DeepHDRVideo [19].
TABLE I: Comparison results on Kalantari’s dataset and Prabhakar’s dataset. The best results are highlighted and the second best results are underlined.
Models 17’CGF 19’CVPR 20’ECCV 22’ECCV 22’MM 23’ICCV 24’ECCV
Kalantari AHDRNet Prabhakar HDR-Trans DomainPlus SCTNet SAFNet Ours
train and test on Kalantari’s dataset
PSNR-µ 42.74 43.77 43.08 44.28 44.02 44.13 44.61 44.69
PSNR-L 41.22 41.35 41.68 42.88 41.28 42.12 43.09 42.19
SSIM-µ 0.9877 0.9907 / 0.9916 0.9910 0.9916 0.9918 0.9920
SSIM-L 0.9848 0.9859 / 0.9884 0.9864 0.9890 0.9892 0.9895
HDR-VDP-2 60.51 62.30 62.21 66.03 62.91 66.65 66.93 68.35
train and test on Prabhakar’s dataset
PSNR-µ 35.63 38.61 38.30 41.31 40.38 41.23 40.18 41.80
PSNR-L 32.50 35.26 34.98 39.44 38.08 38.75 37.90 40.13
SSIM-µ 0.9613 0.9663 0.9702 0.9726 0.9698 0.9724 0.9705 0.9731
SSIM-L 0.9692 0.9794 0.9781 0.9885 0.9872 0.9881 0.9865 0.9895
HDR-VDP-2 59.42 61.14 / 63.01 62.12 62.29 62.04 63.53
TABLE II: Cross-dataset evaluation on Kalantari’s dataset and Prabhakar’s dataset. The best results are highlighted and the second best results are underlined.
2019 2022 2022 2023 2024
Models AHDRNet HDR-Trans DomainPlus SCTNet SAFNet Ours
train on Kalantari’s dataset, test on Prabhakar’s dataset
PSNR-µ 33.96 34.07 32.64 33.83 38.00 39.26
PSNR-L 32.46 36.62 30.4233 30.95 34.65 35.99
SSIM-µ 0.9601 0.9675 0.9046 0.9584 0.9597 0.9707
SSIM-L 0.9542 0.9656 0.9094 0.9521 0.9793 0.9848
train on Prabhakar’s dataset, test on Kalantari’s dataset
PSNR-µ 40.03 41.38 41.15 40.88 40.86 42.02
PSNR-L 36.71 39.21 38.18 37.59 37.50 39.38
SSIM-µ 0.9855 0.9890 0.9873 0.9892 0.9882 0.9903
SSIM-L 0.9758 0.9873 0.9837 0.9842 0.9810 0.9870
Figure 6: Visual comparison on Kalantari’s dataset. While other methods suffer from detail loss, overexposed patches, and color inconsistencies, our approach successfully reconstructs overexposed details and preserves natural, consistent color.
Figure 7: Visual comparison on Prabhakar’s dataset. Most existing methods suffer from ghosting artifacts, our EAFNet mitigates these issues.
Experiments on Prabhakar’s datasets and Kalantari’s datasets.
Intra-Validation. Our DCS can be viewed as a MEF task, so we conduct a comparison with the state-of-the-art MEF method using the publicly available MEF dataset. The quantitative results on both datasets are shown in Table I, our EAFNet achieves superior performance across all datasets. Specifically, on the Kalantari dataset, our method surpasses the second-best approach in terms of PSNR-$\mu$ by 0.08 dB, while on the Prabhakar dataset, it achieves a more substantial improvement of 0.49 dB. The visual comparisons in Fig.  7 illustrate that most existing methods still suffer from noticeable ghosting artifacts. In contrast, the proposed EAFNet effectively mitigates these issues. Moreover, in overexposed regions (Figure 6), we find that the Kalantari method leads to detail loss, while AHDRNet and HDR-Trans exhibit large white overexposed patches in wall areas. Additionally, methods such as DomainPlus still struggle with color inconsistencies. In contrast, our approach successfully reconstructs most of the overexposed details while preserving natural and consistent color reproduction.
Figure 9: Visual comparison on Sen’s dataset. Existing methods suffer from misalignment or detail loss when the reference image is over-exposed, while our method reconstructs clearer and more accurate textures.
Cross-Validation. The cross-testing on the two datasets serves to mitigate any potential bias and offers a comprehensive understanding of the model’s reliability across varying conditions. As shown in Table II, we achieve superior performance in cross-dataset evaluation, demonstrating the robustness and generalizability of our model.
Unsupervised-dataset. To assess the generalization performance of our SAFNet, we conduct qualitative comparisons on the Sen’s and Tursun’s datasets. As shown in Fig. 8, when there exists a significant luminance contrast between the reference and non-reference images, such as the case where white backgrounds in the reference image correspond to dark content in the non-reference frames, most existing methods tend to produce prominent visual artifacts. In contrast, the proposed SAFNet effectively suppresses artifacts, delivering perceptually plausible HDR reconstructions even in scenes with severe exposure disparity. In Fig. 9, When the reference image is over-exposed and misalignment exists between the reference and non-reference images, methods such as AHDRNet, DominPlus, and SAFNet exhibit incorrect alignment, while HDR-Trans and SCTNet suffer from varying degrees of detail loss. In contrast, our method is capable of reconstructing clearer and more accurate textures.
@@ V-A Comparison with state-of-the-art methods
Figure 10: Visual comparison with MEF methods on self-captured videos. Our EAFNet reduces ghosting from motion (first row) and background misalignment (second row), recovering natural colors and fine details in dark regions (third row).
@@
TABLE III: Comparison on self-captured videos. * indicates MEF methods under the AE setup. The reported time refers to the inference time for a 128×128 patch on a single RTX 4090 GPU.
AE setup MEF setup with DCS
21’ICCV 23’ICCV 24’CVPR 23’ICCV 24’ECCV 22’ECCV 22’MM 23’ICCV 24’ECCV
Chen LAN-HDR HDRFlow SCTNet* SAFNet* HDRTrans DomainPlus SCTNet SAFNet Ours
LSD 0.1031 0.1003 0.0914 0.1027 0.0976 0.0093 0.0094 0.0093 0.0097 0.0093
t-SSIM 0.5913 0.6879 0.7352 0.6947 0.6862 0.8923 0.8991 0.8868 0.8851 0.9071
$L_{avg}$ 105.02 108.75 113.49 109.23 111.70 111.24 110.41 106.84 111.32 133.16
Runtime(ms) 22.23 21.69 2.6 11.17 3.2 21.6 7.44 11.17 3.2 4.3
V-B Experiments on self-captured videos
Overview. In this section, we explore the effectiveness of our DCS in real-world scenes and assess the capabilities of the proposed EAFNet in HDR video reconstruction. As detailed in Section III, real video data is captured using dual MV-CS032-10GC industrial cameras. Due to the alternating exposure method’s requirement for continuous alternating exposure frames, direct comparison with our approach under identical settings is not feasible. To ensure a fair comparison, we simultaneously capture two datasets with our system, ensuring minimal scene variation. The first camera captures alternating low, medium, and high exposure sequences, while the second camera continuously captures a medium exposure sequence. The AE method uses the alternating exposure sequence from the first camera as the test dataset, while our dual-camera framework takes low and high-exposure images from the first camera and a continuous medium-exposure sequence from the second. Both cameras have a resolution of $960\times 540$, and we select nine scenes spanning indoor, outdoor, day, and night conditions, each covering low-light, normal-light, and high-light scenarios with challenges such as high contrast, motion, and occlusion. The gain follows the scene variations, using exposure times of (-2, 0, 2 EV).
Compare with MEF methods. We compare our method with the state-of-the-art MEF methods including HDR-Transformer [40], SCTNet [43], DomainPlus [39], and SAFNet [31]. Firstly, we compare the performance of models trained on two different datasets, Prabhakar’s and Kalantari’s. As shown in Fig. 10 (third row), we find that models trained on Kalantari’s dataset exhibit unnatural textures in certain scenes. Therefore, all MEF results are tested using the model trained on Prabhakar’s dataset. During real video capture, complex lighting conditions, camera and object motion, and high-contrast scenes pose significant challenges for video reconstruction. As shown in the first row of Fig. 10, DomainPlus produces motion misalignment artifacts, while SAFNet and STCNet generate unnatural artifacts in some very dark regions. In contrast, our EAFNet effectively eliminates these artifacts, restoring more natural colors and brightness in darker regions, thus providing more detailed information. The second row of the figure demonstrates that for high-contrast scenes (where reference images are overexposed or areas with white backgrounds correspond to low/high exposure images of dark objects), existing methods fail to eliminate misalignment artifacts. However, our EAFNet can produce more natural reconstruction results, delivering a visually coherent output. The third row shows that our EAFNet provides more detail in the dark region, which leads to high luminance in Table III.
Compare with AE methods. Since the AE method and our method require training on different dataset settings, the AE method is trained on its public dataset following the AE training setup, while our method is trained on the publicly available Prabhakar’s dataset. The models are then tested on our dual-camera dataset. We compare the state-of-the-art AE methods including HDRFlow [51], LAN-HDR [50], and DeepHDRVideo [19]. As shown in Fig. 11 and Table III, for real-world test videos captured in our DCS, the luminance of the video reconstructed using the AE method exhibits abrupt changes, leading to noticeable flickering, which makes it impractical for real-world applications. In contrast, our DCS produces temporally consistent results, which suggests that the proposed system can effectively overcome the challenges of traditional AE methods.
Efficiency Analysis. The overall efficiency of HDR video processing is determined by both the acquisition throughput and the inference speed. Our method specifically targets the acquisition bottleneck by introducing an asynchronous dual-camera design, which decouples the capture of alternate exposure frames from the reference stream. This architecture effectively eliminates the latency caused by long high-exposure frames in traditional single-camera systems, enabling real-time HDR frame acquisition. Although our reconstruction model (EAFNet) is not specifically optimized for real-time inference, as shown in Table III, it still exhibits competitive runtime performance.
Dual-Camera Analysis. The use of a dual-camera system inevitably introduces parallax and temporal misalignment due to the baseline separation and potential desynchronization between the two cameras. In our prototype, the baseline is constrained by mechanical limitations but is minimized to the extent possible. Although some parallax is unavoidable, we find that most MEF-based methods in our system demonstrate reasonable robustness to mild parallax. The training data involves dynamic scenes with moving objects and occlusions, the model thus learns to handle motion-induced spatial misalignments, which share similar characteristics with parallax-related shifts in dual-camera setups. In challenging cases, such as those involving high local contrast, misalignment becomes more difficult to resolve. For instance, in the second row of Figure 10, the presence of strong contrast across exposure frames poses a severe test for existing methods. In typical dynamic scenes with such conditions, most existing approaches struggle to produce satisfactory results. In contrast, our proposed EAFNet effectively mitigates artifacts, preserving detail and structure even under these difficult situations. This robustness is achieved through a combination of a pre-alignment module and a reference-guided fusion strategy, which together reduce visible artifacts even in scenes with high contrast or occlusion.
In addition, the system does not rely on precise camera calibration or strict synchronization. The fusion framework is designed to tolerate both geometric misalignment and temporal offset. As shown in Table IV, ablation studies under various calibration conditions and randomized non-reference input orders confirm that the overall reconstruction quality remains stable. Small changes in t-SSIM are mainly due to image cropping rather than alignment failure.
TABLE IV: t-SSIM comparison under different dual-camera deployment settings
t-SSIM MEF setup with DCS
22’ECCV 22’MM 23’ICCV 24’ECCV
HDRTrans DomainPlus SCTNet SAFNet Ours
ori 0.8923 0.8991 0.8868 0.8851 0.9071
w/o calibration 0.8922 0.8994 0.8872 0.8861 0.9079
random non-reference 0.8926 0.8997 0.8842 0.8848 0.9072
lower frame rate 0.8924 0.8985 0.8871 0.8836 0.9066

### v2

@@ IV-E Loss Function
$\mathcal{A}\mathcal{S}\mathcal{L}(\widehat{Z},Z)=\frac{1}{N}\sum\mid\operatorname{Sobel}^{\star}(\widehat{Z})-\text{Sobel}^{\star}(Z)\mid\\$ (8)
@@
$\mathcal{L}_{total}=\mathcal{L}1(\widehat{Z},Z)+\lambda\cdot\mathcal{D}-\mathcal{AS}\mathcal{L}^{D}(\widehat{Z},Z)$ (10)
@@
TABLE I: Comparison results on Kalantari’s dataset and Prabhakar’s dataset. Throughout this paper, the best and second-best results of each case are highlighted in bold red and underlined blue, respectively.
dataset setting train and test on Kalantari’s dataset train and test on Prabhakar’s dataset
Method metrics PSNR-$\mu$ ($\uparrow$) PSNR-L ($\uparrow$) SSIM-$\mu$ ($\uparrow$) SSIM-L ($\uparrow$) HDR-VDP-2($\uparrow$) PSNR-$\mu$ ($\uparrow$) PSNR-L ($\uparrow$) SSIM-$\mu$ ($\uparrow$) SSIM-L ($\uparrow$) HDR-VDP-2 ($\uparrow$)
Kalantari(CGF 2017) [23] 42.74 41.22 0.9877 0.9848 60.51 35.63 32.50 0.09613 0.9692 59.42
AHDRNet(CVPR 2019) [56] 43.77 41.35 0.9907 0.9859 62.30 38.61 35.26 0.9663 0.9794 61.14
Prabhakar(ECCV 2020) [43] 43.08 41.68 - - 62.21 38.30 34.98 0.9702 0.9781 -
HDR-Trans(ECCV 2022) [34] 44.28 42.88 0.9916 0.9884 66.03 41.31 39.44 0.9726 0.9885 63.01
DomainPlus(MM 2022) [62] 44.02 41.28 0.9910 0.9864 62.91 40.38 38.08 0.9698 0.9872 62.12
SCTNet(ICCV 2023) [46] 44.13 42.12 0.9916 0.9890 66.65 41.23 38.75 0.9724 0.9881 62.29
SAFNet(ECCV 2024) [27] 44.61 43.09 0.9918 0.9892 66.93 40.18 37.90 0.9705 0.9865 62.04
EAFNet(Ours) 44.69 42.19 0.9920 0.9895 68.35 41.80 40.13 0.9731 0.9895 63.53
TABLE II: Cross-dataset evaluation on Kalantari’s dataset and Prabhakar’s dataset.
dataset setting train on Kalantari’s dataset, test on Prabhakar’s dataset train on Prabhakar’s dataset, test on Kalantari’s dataset
Method metrics PSNR-$\mu$ ($\uparrow$) PSNR-L ($\uparrow$) SSIM-$\mu$ ($\uparrow$) SSIM-L ($\uparrow$) PSNR-$\mu$ ($\uparrow$) PSNR-L ($\uparrow$) SSIM-$\mu$ ($\uparrow$) SSIM-L ($\uparrow$)
AHDRNet(CVPR 2019) [56] 33.96 32.46 0.9601 0.9542 40.03 36.71 0.9855 0.9758
HDR-Trans(ECCV 2022) [34] 34.07 36.62 0.9675 0.9656 41.38 39.21 0.9890 0.9873
DomainPlus(MM 2022) [62] 32.64 30.42 0.9046 0.9074 41.15 38.18 0.9873 0.9837
SCTNet(ICCV 2023) [46] 33.83 30.95 0.9584 0.9521 40.88 37.59 0.9892 0.9842
SAFNet(ECCV 2024) [27] 38.00 34.65 0.9597 0.9793 40.86 37.50 0.9882 0.9810
EAFNet(Ours) 39.26 35.99 0.9707 0.9848 42.02 39.38 0.9903 0.9870
@@ V EXPERIMENTS
In this section, we conduct extensive experiments to evaluate the performance of our proposed dual-stream paradigm in HDR video generation. Our dual-stream paradigm, supported by a dual-camera system and the EAFNet model, aims to address key challenges in HDR video reconstruction, such as temporal luminance consistency and spatial deghosting.
First, in Section V-A, we outline the experimental setup, including details on the training process, datasets, baseline methods, and evaluation metrics. In Section V-B, we focus on dynamic HDR image fusion, presenting both quantitative and qualitative comparisons with state-of-the-art HDR image deghosting methods to validate the effectiveness of our method in single-frame spatial fusion and deghosting. Section V-C shifts to HDR video reconstruction, where we assess the advantages of our paradigm in temporal consistency and flicker suppression, comparing it with existing AE paradigms and further validating the benefits of the EAFNet model within our proposed paradigm. Section V-D presents an ablation study, analyzing the impact of the dual-camera system configuration and the contributions of each EAFNet component. Finally, Section V-E discusses the limitations of the proposed paradigm, including the effects of camera parallax and system efficiency, and highlights potential directions for future improvements.
V-A Experiment Settings
@@
Training Details. The proposed method is implemented in PyTorch on an NVIDIA RTX4090 GPU. For the convolutional layers, we employ 64 kernels of size $\times3$ with a stride of one, applying zero padding to preserve the dimensions of the resulting feature maps. The optimization is performed using the Adam optimizer [26] with an initial learning rate of $10^{-4}$, and the training concludes when the learning rate reaches $10^{-6}$. To train the model, both the LDR and corresponding images are cropped into 256 $\times$ 256 patches. A batch size of 16 is utilized during training. Patches are randomly rotated for data augmentation to avoid overfitting during the training stage.
Datasets. High-quality HDR video datasets suitable for supervised training or frame-wise quantitative evaluation are scarce, as most lack per-frame HDR ground truth and sufficient exposure diversity. To fairly and comprehensively evaluate our method, we separate the evaluation into two complementary parts: (1) Spatial evaluation focuses on single-frame fusion accuracy using HDR image datasets with reliable ground truth, enabling objective comparison in reconstruction quality. We assess single-frame fusion quality using three widely used HDR image datasets: Prabhakar’s dataset [44], Kalantari’s dataset [22], and Tursun’s dataset [48],
@@
  • Prabhakar’s Dataset. Including 466 training samples and 116 testing samples, with exposure values between -3EV and +3EV.
@@
  • Kalantari’s dataset. Including 74 training samples and 15 testing samples. Each consists of three LDR images, with exposure values of {-2 EV, 0 EV, +2 EV} or {-3 EV, 0 EV, +3 EV}.
@@
  • Tursun’s dataset. Lacking ground truth, is used to assess the generalization ability of the models.
(2) Temporal evaluation focuses on luminance consistency and flicker suppression in dynamic scenes, which cannot be assessed using static image datasets. To enable a comprehensive comparison of our dual-stream HDR video generation paradigm against the widely adopted AE paradigm, we construct dedicated datasets tailored to each paradigm under real-world scenes using our DCS. This approach is necessary because the two paradigms differ fundamentally in their acquisition strategies; a direct comparison using identical inputs would be inherently unfair.
Specifically, we capture two synchronized video streams for each scene using two identical cameras (resolution $960\times 540$). This synchronization is applied only during data acquisition to minimize scene variations between paradigms. In practical deployment, our dual-camera system operates asynchronously without hardware-level synchronization (see Section III). The first camera follows an alternating exposure pattern, and the second camera continuously captures medium-exposure frames. Our paradigm uses the low- and high-exposure frames from the first camera together with the medium-exposure frames from the second camera to reconstruct HDR video. For the AE pipeline, only the alternating exposure sequence from the first camera is used, emulating the standard AE pipeline.
We capture 26 real-world scenes under matched conditions for both datasets, covering diverse scenarios including indoor/outdoor environments, day/night lighting, and challenging situations such as low-light, high-contrast scenes, object motion, and occlusions. Exposure settings are fixed to (2,0,+2)(-2,0,+2) or (1,0,+1)(-1,0,+1) EV to ensure sufficient exposure diversity across varying lighting conditions. This paired capture protocol guarantees that both paradigms are evaluated on scene-matched content while preserving their respective optimal acquisition strategies. However, since the dataset is captured in real-world scenes, there are two challenges in ensuring a fair comparison. First, slight parallax exists between the two camera streams, as the cameras are not perfectly aligned. To mitigate this, we capture predominantly long-range scenes, which minimize the impact of parallax. Second, temporal asynchrony arises from the exposure duration differences inherent in the AE settings. This asynchrony is a fundamental aspect of the AE paradigm and cannot be avoided. With these considerations, our paired capture protocol ensures that both paradigms are evaluated on scene-matched content while preserving their respective optimal acquisition strategies.
Figure 6: Visual comparison on Kalantari’s and Prabhakar’s dataset. While other methods suffer from detail loss, overexposed patches, and color inconsistencies, our approach successfully reconstructs overexposed details and preserves natural, consistent color.
Baselines. We quantitatively and qualitatively compare our EAFNet to several state-of-the-art HDR image deghosting methods including Kalantari [23], AHDR [56], Prabhakar [43], HDR-Transformer [34], SCTNet [46], DomainPlus [62], and SAFNet [27]. For fairness, we exclude few-shot and unsupervised methods that are not designed for supervised multi-exposure HDR reconstruction.
We further compare our proposed framework with representative AE-based HDR video reconstruction methods, including HDRFlow [54], LAN-HDR [8], and DeepHDRVideo [2]. Due to differences in system assumptions and training strategies, AE methods are trained on their respective public datasets following the official configurations, while our method is trained on the publicly available Prabhakar dataset. All models are evaluated on the same real-world video sequences captured using our dual-camera system.
@@
Metrics. We quantitatively evaluate the predicted HDR images by measuring PSNR and SSIM [52] in both the linear domain (-L) and the tone-mapped domain (-$\mu$) [22]. The tone mapping operation is described in Sec. IV-E. Additionally, we compare the results using HDR-VDP-2 [37] for conforming the human visual perception. We also quantitatively evaluated the luminance consistency of the video by testing the Luminance Standard Deviation (LSD), temporal SSIM (t-SSIM) and average luminance ($L_{avg}$):
@@
TABLE III: Comparison on self-captured videos. * indicates HDR image deghosting methods under the alternating exposure solution. The reported time refers to the inference time for a 128×128 patch on a single RTX 4090 GPU.
Method Paradim LSD ($\downarrow$) t-SSIM ($\uparrow$) $L_{avg}$ Inference time (ms, $\downarrow$)
DeepHDRVideo(ICCV 2021) [2] 0.1031 0.5913 105.02 22.23
LAN-HDR(ICCV 2023) [8] 0.1003 0.6879 108.75 21.69
HDRFlow(CVPR 2024) [54] 0.0914 0.7352 113.49 2.62
SCTNet*(ICCV 2023) [46] 0.1027 0.6947 109.23 11.17
SAFNet*(ECCV 2024) [27] alternating-exposure paradigm 0.0976 0.6862 111.70 3.27
HDR-Trans(ECCV 2022) [34] 0.0093 0.8923 111.24 21.60
DomainPlus(MM 2022) [62] 0.0094 0.8991 110.41 7.44
SCTNet(ICCV 2023) [46] 0.0093 0.8868 106.84 11.17
SAFNet(ECCV 2024) [27] our dual-stream paradigm 0.0097 0.8851 111.32 3.24
EAFNet(Ours) 0.0093 0.9071 133.16 4.32
V-B HDR Image Experimental Results
Quantitative Comparisons. Table I reports the intra-dataset evaluation results on Prabhakar’s and Kalantari’s datasets. Our EAFNet consistently achieves the best performance across both datasets. On the Kalantari dataset, it surpasses the second-best method by $0.08$ dB in PSNR-$\mu$, while on the Prabhakar dataset, the margin increases to $0.49$ dB. These gains are complemented by improvements in SSIM-$\mu$, indicating that our exposure-adaptive fusion strategy benefits both fidelity and structural consistency.
We further conduct cross-dataset validation, where training and testing are performed on different datasets (Table II). This setting is critical for evaluating the robustness of HDR fusion models against variations in exposure patterns, motion distributions, and scene content, thereby better reflecting real-world deployment scenarios. Our EAFNet maintains clear superiority in this challenging setting, with cross-domain gains even larger than in the intra-dataset case. For example, when trained on Kalantari and tested on Prabhakar, EAFNet achieves 39.26 dB PSNR-$\mu$ and 0.9707 SSIM-$\mu$, surpassing the next best approach by over 1.2 dB and 0.03, respectively. This demonstrates that our model does not overfit to dataset-specific statistics, but learns exposure-aware and motion-robust fusion representations that transfer effectively across domains. The strong bidirectional results confirm the generality and domain-agnostic nature of our fusion mechanism.
Qualitative Comparisons. Fig.6 presents visual comparisons on the Prabhakar and Kalantari datasets. Most existing methods exhibit noticeable ghosting artifacts, particularly around moving objects and edges with exposure differences. In overexposed regions, the Kalantari method suffers from detail loss, while AHDRNet and HDR-Trans produce large saturated patches, and DomainPlus struggles with color inconsistencies. In contrast, EAFNet recovers fine structures in saturated areas while preserving natural and consistent color reproduction.
We also evaluate on challenging HDR datasets without ground truth. In Tursun’s dataset (Fig.7), large luminance differences between reference and non-reference frames often cause severe artifacts in competing methods, particularly when bright backgrounds in the reference correspond to dark content in non-reference frames. EAFNet effectively suppresses these artifacts, producing perceptually coherent HDR reconstructions.
@@ V-B HDR Image Experimental Results
Figure 8: Visual comparison with HDR image deghosting methods on self-captured videos. Our EAFNet reduces ghosting from motion (first row) and background misalignment (second row), recovering natural colors and fine details in dark regions (third row).
@@
V-C HDR Video Experimental Results
Quantitative Comparisons. Table III summarizes the quantitative results. We interpret these findings from two perspectives:
Paradigm-level comparison (AE vs. our dual-stream). Dual-stream-based methods consistently achieve lower luminance standard deviation (LSD) and higher temporal SSIM (t-SSIM) than AE-based methods, indicating superior temporal stability and reduced flickering. For instance, the best AE-based method, HDRFlow, attains an LSD of 0.0914 and a t-SSIM of 0.7352, whereas our EAFNet under the dual-stream paradigm achieves 0.0093 and 0.9071, respectively. This substantial improvement stems from the fixed-exposure temporal reference in the dual-stream pipeline, which effectively eliminates inter-frame luminance flickering caused by alternating exposure in AE pipelines.
Within-dual-stream comparison (our EAFNet vs. image deghosting baselines). When all methods operate under our dual-stream paradigm, our EAFNet outperforms all integrated image deghosting-based baselines in both temporal and spatial quality metrics. EAFNet achieves the highest t-SSIM (0.9071) and ties for the best LSD (0.0093), reflecting its ability to simultaneously preserve temporal coherence and fine-grained spatial details. Notably, it accomplishes this while maintaining a reasonable inference time (4.32 ms for a $128\times 128$ patch), which is faster than most transformer-based baselines.
Qualitative Comparisons. Visual comparisons in Fig. 9 and Fig. 8 further support the quantitative findings.
Paradigm-level comparison (AE vs. our dual-stream). As shown in Fig. 9, AE-based reconstructions exhibit noticeable luminance fluctuations between consecutive frames. These artifacts are inherent to AE pipelines due to the lack of a consistent exposure reference. In contrast, our results maintain stable luminance across frames and significantly reduce temporal artifacts, benefiting from the persistent medium-exposure reference stream.
Within-dual-stream comparison (our EAFNet vs. image deghosting baselines). As shown in Fig. 8, in high-motion scenarios, some image deghosting baselines (e.g., DomainPlus, SAFNet) generate ghosting or introduce color distortions in dark regions, whereas our EAFNet preserves spatial alignment and reproduces accurate color. Under extreme luminance contrast, EAFNet better suppresses misalignment artifacts, maintaining natural transitions in overexposed or underexposed areas. In localized low-light regions, EAFNet recovers fine structures without introducing noise, enhancing perceptual clarity. These visual improvements align with the quantitative margins in Table III, highlighting the advantage of EAFNet within the dual-stream framework.
V-D Ablation study
In this section, we first evaluate the robustness of our framework under different dual-camera deployment conditions using our self-captured real-world dataset, focusing on the impact of calibration accuracy, input ordering, and frame rate. We then perform an ablation study on EAFNet to analyze the contribution of each component and the source of efficiency, with all quantitative results reported on Prabhakar’s dataset [44].
Dual-Camera Analysis. In a dual-camera system, imperfect calibration and potential desynchronization can lead to geometric misalignment and motion-induced misalignment, respectively. In our prototype, no special effort is made to achieve sub-pixel calibration or perfect synchronization. Instead, the fusion framework is designed to tolerate moderate deviations in both geometry and timing. The training data include dynamic scenes with motion and occlusion, allowing the model to learn robustness to such spatial inconsistencies.
Table IV reports t-SSIM under four deployment variants. Variations across deployment variants are small for all methods ($\leq$0.003), indicating that, within the tested range, moderate calibration errors, input order changes, and frame-rate reductions have limited impact on temporal flickering. Minor fluctuations are likely due in part to the cropping required for alignment rather than alignment failure itself. Overall, the results suggest that, within the tested range of deviations, temporal stability can be maintained without the need for precise calibration and strict synchronization, but a higher level of calibration and synchronization may still be beneficial for optimizing performance in more demanding scenarios.
TABLE IV: t-SSIM comparison under different dual-camera deployment settings. base: Baseline dual-camera system settings. w/o Cal.: Dual-camera system without calibration. Random nr.: Randomized input order of non-reference frames. Lower fr.: Lower frame rate for non-reference frame input.
Method settings Base. w/o Cal. Random nr. Lower fr.
HDR-Trans(ECCV 2022) [34] 0.8923 0.8922 0.8926 0.8924
DomainPlus(MM 2022) [62] 0.8991 0.8994 0.8997 0.8985
SCTNet(ICCV 2023) [46] 0.8868 0.8872 0.8842 0.8871
SAFNet(ECCV 2024) [27] 0.8851 0.8861 0.8848 0.8836
EAFNet(Ours) 0.9071 0.9079 0.9072 0.9066